For what value of x will the line passing through the points A(x+14, 5) and B(2−3x, 8) be vertical?

Answer by ikleyn(52946)

.

It will be vertical, when the first coordinates (x-coordinates) of both given points are the same


    x+14 = 2-3x,


or


    x + 3x = 2 - 14

    4x = -12

      x = -12/4 = -3.    ANSWER